# Resonant Solitary States in Complex Networks

### v1

@@ Abstract
$m_{z}\ddot{\vartheta}_{z}=$ $P_{z}-\alpha_{z}(\dot{\vartheta}_{z}+\omega_{s})-p_{s}-p_{zk}^{\text{osc}}(t),$ (7)
where $L^{*}$ is the effective coupling Laplacian on the synchronized cluster with weights $\kappa_{ij}\cos(\vartheta^{*}_{i}-\vartheta^{*}_{j})$.
Note that $p_{s}$ does not appear in the equation for $S$, since the synchronized cluster steady state accounts for the average power injection from the solitary. We also note that the dependence of $L^{*}$ on $p_{s}$ is typically weak (SM I.B).
For $\vartheta_{z}$ to stay bounded, and for $\omega_{s}$ to be the long-term average of the solitary’s frequency, the long-term average $\vartheta_{z}$ has to be zero. This provides us with the condition $0=P_{z}-\alpha_{z}\omega_{s}-p_{s}$. $p_{s}$ is a measure for the mean coupling of $S$ and $z$, while $\omega_{s}$ stands for their incoherence. Note that this relationship between $\omega_{s}$ and $p_{s}$ can be observed in Fig. 1. In synchrony there is a steady high energy flow. In the natural solitary state, the mean energy flow is zero and the system is effectively decoupled. For intermediate solitary states, there is some intermediate amount of energy.
If we now can express $p_{s}$ as a function of $\omega_{s}$, this equation becomes a self-consistent equation for the frequency of the solitary state:
Additionally, a potentially stable solitary state requires the correct sign of the change in $Z$ with frequency. If we imagine $\omega_{s}$ as slowly varying, we have $m_{z}\dot{\omega}_{s}=Z(\omega_{s})$. Thus, we expect the sign of the derivative $\partial_{\omega_{s}}Z(\omega_{s})$ to be informative about the stability of the solitary state, at least heuristically.
Both $p_{s}$ and $\omega_{s}$ depend on the trajectory of the solution. To obtain the relationship between them, we will leverage averaging theorems Guckenheimer and Holmes (2002) to obtain an approximate solution for which they can be exactly evaluated. To do so, we introduce a timescale $\varepsilon^{-1}$ that splits the slow, non-rotating and the rotating dynamics. By rescaling the time axis we can eliminate $m$ in Eq. (1), and set $\alpha$, $P$, and $\kappa$ to $\varepsilon$, $\varepsilon P$ and $\varepsilon\kappa$ for a small value $\varepsilon$. We choose $\varepsilon$ such that while the rescaled $\kappa$ remains of order one, the eigenvalues of the rescaled effective Laplacian $L^{*}$ for the eigenmodes localized at $k$ are of order $\varepsilon^{-1}$. As a result, we treat $L^{r}=\varepsilon L^{*}$ as being of order one. Such a scaling can be achieved, if node $k$ has a sufficiently high degree (SM I.A).
After this rescaling, Eq. (6) and Eq. (7) become
@@
Equation (9) has the form of the perturbation of the Hamiltonian system $\ddot{\vec{\vartheta}}=-L^{r}\vec{\vartheta}$, where the perturbation is due to $\vartheta_{z}$. The appropriate perturbative approach in this case is averaging Guckenheimer and Holmes (2002), where one should first write the system for the slowly varying amplitudes of the network modes, i.e., the amplitudes of the periodic solutions of the unperturbed system, see Eq. (S48). The amplitude dynamics is then decomposed into modes resonant with $\omega_{c}$, the driving frequency in $p_{zk}^{\text{osc}}$, and non-resonant ones. The contributions of both types of modes are small, allowing us to average the dynamics of $\vec{\vartheta}$. We obtain $\vartheta_{k}(t)$ and insert it into Eq. (10). We then average Eq. (10) and undo the rescaling in this equation to obtain the explicit form of the self-consistent Eq. (8):
$Z(\omega_{s})=P_{z}-\alpha_{z}\omega_{s}-\frac{\kappa_{zk}^{2}}{2}\sum_{\ell=1}^{N-1}\frac{\alpha\omega_{c}\left(v_{k}^{[\ell]}\right)^{2}}{\left(\lambda^{[\ell]}-m\omega_{c}^{2}\right)^{2}+\alpha^{2}\omega_{c}^{2}}.$ (11)
@@
In Eq. (11), the sum runs over eigenmodes $\ell$ of $L^{*}$, with eigenvalue $\lambda^{[\ell]}$ and eigenvector $v^{[\ell]}$. To efficiently evaluate Eq. (11), we assume that the dependence of $L^{*}$ and its eigenmodes on $p_{s}$ is weak. This is a reasonable assumption if the synchronized cluster is significantly larger than the solitary cluster (node $k$) and the energy injected at node $k$ dissipates quickly into $S$.
@@
A linear response picture.— The obtained form of $p_{s}$ has a natural interpretation. The solitary state, with frequency $\omega_{c}$, perturbs the synchronized cluster. The response of the cluster has the frequency $\omega_{c}$ and some well-defined phase shift. This enables a non-zero energy flow $p_{s}$ between $S$ and $z$, which is responsible for the frequency shift $\omega_{s}$. The flow is proportional to the amplitude of the response and the local amplitude of the corresponding eigenvector. Thus, large energy flows can occur if $\omega_{c}$ is close to resonance with an eigenmode of $S$ that is localized at $k$. As $\omega_{c}$ shifts with the energy flow, the system can robustly tune itself towards such resonances if they are present. This resonant tuning mechanism predicts why certain topological features, namely the ones causing highly localized modes, enable the generation of additional solitary states.
@@
Following this interpretation, we can directly derive the result for $Z(\omega_{s})$ in Eq. (11) using linear response theory, if we neglect the nonlinear part of $p_{zk}^{\text{osc}}$ in Eq. (9), which is eventually averaged out in the averaging approach, from the outset. This alternative assumption can be justified, when this nonlinearity is only one of several contributions to the overall coupling of a well-connected node $k$. In this case, the network is harmonically driven by $\sin\omega_{c}t$, and linear response similar to Zhang et al. (2019) can be used to calculate its response, including the amplitudes $a_{k}^{(k)[\ell]}(\omega_{c})$ and phase lags $\delta^{[\ell]}(\omega_{c})$ at the driven node $k$. These then provide us with a time averaged energy flow:
$p_{s}(\omega_{c})=-\frac{\kappa_{kz}}{2}\sum_{\ell=1}^{N}a_{k}^{(k)[\ell]}(\omega_{c})\sin\delta^{[\ell]}(\omega_{c}).$ (12)
@@
Both approaches used to derive Eq. (11) straightforwardly generalize to a solitary node connected to several nodes in $S$. The back-reactions of the network combine linearly, and we simply sum over all neighbors of $z$ (Eq. (S71)).
@@
Examples.— We illustrate the power of Equation (11) using two examples. First we give a minimal effective model that exhibits the resonant tuning mechanism cleanly; then a full complex network with a complex resonant response. We compare the results with numerical simulations Plietzsch et al. (2022). Further examples can be found in the SM II.B.
@@
Figure 3: Network modes shape the landscape of solitary attractor states for the minimal effective model (inset). (left) Intersections between the straight line $P_{1}-\alpha\omega_{s}$ and the mean energy flow $p_{s}(n=6)$ generate stable and unstable solutions according to Eq. (11). We use $P_{1}=1$, $\alpha=0.1$. The single peak of $p_{s}$ is centered around the network mode close to $\sqrt{Kn}$. (right) Bifurcation diagram with the bifurcation parameter $n$. The two stable branches are connected by an unstable branch and centered around the network mode.
@@
The simplest model that shows tristability with two distinct solitary frequencies is a system that neglects the response of the network beyond the node $k$. We have a solitary node ($z=1$), its neighbor ($k=2$), and the rest of the network is assumed to have infinite inertia, and thus no dynamics. This model is shown in the inset in Fig. 3 (right). The neighbor has $n$ links into the infinitely heavy part, each with coupling strength $K$. See SM II.A for details.
@@
The left panel in Fig. 3 shows the energy absorbed by frequency adaptation, a straight line given by $P_{1}-\alpha\omega_{s}$, and the energy flow due to the network’s response, $p_{s}=\langle p_{12}\rangle$. There is a single peak centered around the only network mode, which is close to $\lambda=\sqrt{Kn}$. At the intersection of these two curves, we have a zero of Eq. (11). However, the linear stability heuristic suggests that only intersections with $\partial_{\omega_{s}}Z<0$ are candidates for stable solitary states. Graphically, these correspond to intersections where the network response intersects from below when reading left to right.
@@
Hence, for $n\lesssim 12$, there is a pair of a stable and an unstable solution due to the peak, and a third, stable solution at $\omega_{s}\approx\Omega_{1}=P_{1}/\alpha$. For larger values of $n$, the second and third solutions annihilate, while the first converges to $\omega_{s}=\Omega_{1}$. The bifurcation diagram in the right panel of Fig. 3 illustrates this behavior. Numerical simulations with random initial conditions closely correspond to the predicted frequency and stability of solitary states, confirming the existence of tristability between the synchronous state and solitary states at multiple frequencies, and the overall accuracy of our results.

### v3

@@ IV Averaging
$m_{z}\ddot{\vartheta}_{z}=$ $P_{z}-\alpha_{z}(\dot{\vartheta}_{z}+\omega_{\text{sync}})-p_{s}-p_{zk}^{\text{osc}}(t),$ (14)
where $L^{*}$ is the effective coupling Laplacian on the synchronized cluster with weights $L^{*}_{ij}=\kappa_{ij}\cos(\vartheta^{*}_{i}-\vartheta^{*}_{j})$ and eigenvectors $\vec{v}^{[\ell]}$. We have successfully transformed into a frame that accounts for the power imbalance due to $p_{s}$. This can be seen from the fact that $p_{s}$ does not appear in the equation for $S$, Eq. (13), since the synchronized cluster steady state accounts for the average power injection $p_{s}$ from the solitary node with the $\omega_{\text{sync}}$ rotation.
Second, we need to properly take care of the fact that Eqs. (13-14) become a non-autonomous system when we insert the explicitly time-dependent ansatz Eqs. (6-7) into $p_{zk}$. We introduce a timescale $\varepsilon^{-1}$ that splits the autonomous and non-autonomous parts in Eqs. (13-14). It allows us to write the system in a perturbative form, in which averaging can be applied. Such a scaling can be achieved, if node $k$ has a sufficiently high degree (SM I.A). In that case, there is a Laplacian eigenvalue $\lambda^{[r]}$ that is highly correlated to the node’s degree $d_{k}$. We will observe in the examples below that the solitary frequency tunes itself towards resonance with such eigenmodes: $\omega_{s}^{2}\sim\lambda^{[r]}$.
The rescaling is given by
$t=\sqrt{\frac{m}{\lambda^{[r]}}}t^{\prime},\qquad P_{i}=\frac{P^{\prime}_{i}\alpha\sqrt{\lambda^{[r]}}}{\sqrt{m}},\qquad\kappa_{ij}=\frac{\kappa^{\prime}_{ij}\alpha\sqrt{\lambda^{[r]}}}{\sqrt{m}},$ (15)
dropping the prime from now on. Further, we set $\varepsilon=\alpha/\sqrt{m\lambda^{[r]}}\ll 1$. In other words, by rescaling the time axis we can absorb $m$ in Eq. (13), and set $\alpha$, $P$, and $\kappa$ to $\varepsilon$, $\varepsilon P$ and $\varepsilon\kappa$. We choose $\varepsilon$ such that, while the rescaled $\kappa$ remains of order one, the eigenvalues of the rescaled effective Laplacian $L^{*}$ for the eigenmodes localized at $k$ are of order $\varepsilon^{-1}$. As a result, we treat $L^{r}=\varepsilon L^{*}$ as being of order one. Eqs. (13-14) become
@@
Third, a crucial step for solvability is to approximate $\varepsilon p_{zk}(t)$, cf. Eq. (12), to first order in $\varepsilon$, which finally allows us to identify an explicit expression for $p_{s}$ in terms of $\vartheta_{k}$ and $\omega_{c}$,
$\varepsilon p_{zk}(t)$ $\approx\varepsilon\kappa_{zk}\left(\sin\omega_{c}t-\vartheta_{k}(t)\cos\omega_{c}t\right),$ (18)
$p_{s}$ $\approx\kappa_{zk}\langle\vartheta_{k}(t)\sin\omega_{c}t\rangle.$ (19)
@@
Equation (16) has the form of the perturbation of the Hamiltonian system $\ddot{\vec{\vartheta}}=-L^{r}\vec{\vartheta}$, where the perturbation is of order $\varepsilon$. The perturbation is caused partly by the presence of $\vartheta_{z}(t)$, and by other small terms, such as damping. The appropriate perturbative approach in this case is averaging Guckenheimer and Holmes (2002), where one should first write the system for the slowly varying amplitudes $\vec{x}(t)$ and $\vec{z}(t)$ of the network modes $\vec{\xi}(t)$, i.e., the amplitudes of the periodic solutions of the unperturbed system. This is achieved by diagonalization of the Laplacian with the linear transformation $Q_{ij}\coloneqq v_{i}^{[j]}$,
$\vec{\xi}(t)\coloneqq Q^{T}\vec{\vartheta}(t)\coloneqq\vec{x}(t)\cos\omega_{c}t-\vec{z}(t)\sin\omega_{c}t.$ (20)
The amplitude dynamics for $\vec{x}(t)$ and $\vec{z}(t)$ can be inferred with the invertible van der Pol transformation (spelled out in Eq. (S50)) into the frame rotating at the driving frequency $\omega_{c}$. These amplitude dynamics is then decomposed into modes resonant with $\omega_{c}$ and non-resonant modes. For the sake of brevity and clarity, we refer to SM I for details. Most importantly, the contributions of both types of modes are small for different reasons. On one hand, the resonant modes give small contributions due to a small prefactor $\varepsilon\Delta\sim\omega_{c}^{2}-\lambda^{[r]}$. On the other hand, the non-resonant modes, where $\varepsilon\Delta$ is of $O(1)$, give small contributions due to their weak localization, allowing us to average the dynamics of the slowly varying $\vec{x}(t)$ and $\vec{z}(t)$. We obtain an autonomous system that is a good approximation ($\varepsilon$-close) of the non-autonomous system for time scales up to $\varepsilon^{-1}$ Guckenheimer and Holmes (2002). This averaged system (Eq. (S55)) is linear with a unique solution that gives us the average amplitudes $\vec{x}^{*}$ and $\vec{z}^{*}$ of all modes. From those, we obtain the average trajectory of $\vartheta_{k}(t)$, and insert it into Eq. (19) to get an expression for $p_{s}(\omega_{s})$. This finally provides us with the explicit form of the self-consistent Eq. (8):
$0=Z(\omega_{s})$ $=P_{z}-\alpha_{z}\omega_{s}-p_{s}$ (21)
$=P_{z}-\alpha_{z}\omega_{s}-\frac{\kappa_{zk}^{2}}{2}\sum_{\ell=1}^{N-1}\frac{\alpha\omega_{c}\left(v_{k}^{[\ell]}\right)^{2}}{\left(\lambda^{[\ell]}-m\omega_{c}^{2}\right)^{2}+\alpha^{2}\omega_{c}^{2}}.$
@@
In Eq. (21), the sum runs over eigenmodes $\ell$ of $L^{*}$, with eigenvalue $\lambda^{[\ell]}$ and eigenvector $\vec{v}^{[\ell]}$. To efficiently evaluate Eq. (21), we assume that the dependence of $L^{*}$ and its eigenmodes on $p_{s}$ is weak. This is a reasonable assumption if the synchronized cluster is significantly larger than the solitary cluster (node $z$), and the energy injected at node $k$ dissipates quickly into $S$. Practically, the $\vartheta_{i}^{*}$, that depend on $p_{s}$, can be well approximated by the $\varphi_{i}^{*}$, that do not (SM I.B).
Lastly, to close Eq. (21), we need to relate the solitary frequency $\omega_{s}$ to $\omega_{c}$, its relative value to $S$. It can be shown by summing over Eq. (1) for all $i\in S$ and $i=z$, that the $\alpha$-weighted sum of frequencies decays towards zero from all initial conditions, hence $(N-1)\alpha\omega_{\text{sync}}+\alpha_{z}\omega_{s}=0$ (see SM I.A.). This gives us
This linear relationship between $\omega_{s}$ and $\omega_{c}$ is useful in solving Eq. (21).
Having obtained an explicit, closed, self-consistent equation for $\omega_{s}$ from our ansatz, we now turn to its interpretation and evaluation to make predictions of solitary attractor states and compare them to simulations.
V Interpretation: a linear response picture
In this section, we provide an alternative derivation that is a shortcut to Eq. (21), using linear response theory. For an extensive account of linear response in inertial Kuramoto networks, we refer to Zhang et al. (2019). As a motivation, we start by discussing the natural interpretation that the linear response offers, and that can also be deducted from the resulting Eq. (21).
@@ V Interpretation: a linear response picture
The obtained form of $p_{s}$ in Eq. (21) has a natural interpretation. The solitary state, with frequency $\omega_{s}$, perturbs the synchronized cluster $S$, with relative frequency $\omega_{c}$. The response $\vec{\vartheta}(t)$ of the cluster $S$ oscillates at frequency $\omega_{c}$ with amplitudes $a_{i}$ and some well-defined phase shifts $\delta_{i}$, e.g., $\vartheta_{k}(t)\sim a_{k}\sin\left(\omega_{c}t+\delta_{k}\right)$. The phase shift $\delta_{k}$ enables a non-zero energy flow $p_{s}$ between $S$ and $z$, which is responsible for maintaining the relative frequency shift $\omega_{c}$. The energy flow is proportional to the amplitude of the response, and especially the local amplitude of the corresponding eigenvector $v_{k}^{[\ell]}$ for each mode $\ell$. Thus, large energy flows can occur if $\omega_{c}$ is close to resonance with an eigenmode of $S$ that is localized at $k$. As $\omega_{c}$ determines the magnitude of energy flow, but also shifts with the energy flow itself, the system can robustly tune itself towards such resonances if they are present. This resonant tuning mechanism predicts why certain topological features, namely the ones causing highly localized modes, enable the generation of additional solitary states, cf. section VI.
@@
Following this interpretation, we can directly derive the result for $Z(\omega_{s})$ in Eq. (21) using linear response theory, see SM I.C. We observe that the nonlinear part of $p_{zk}^{\text{osc}}(t)$ in Eq. (16) is eventually averaged out in the averaging approach. Linear response can be readily applied, if we neglect this nonlinearity from the outset. The assumption that this nonlinearity is sufficiently small can be justified, when it is only one of several contributions to the overall coupling of a well-connected node $k$. In this case, the cluster $S$ is harmonically driven by $\sin\omega_{c}t$, and linear response similar to Zhang et al. (2019) can be used to calculate its response for each network mode $\ell$ separately, including the modal amplitudes $a_{i}^{(k)[\ell]}(\omega_{c})$ at node $i$ and modal phase lags $\delta^{[\ell]}(\omega_{c})$ given by
$a_{i}^{(k)[\ell]}(\omega_{c})$$\coloneqq$ $\frac{\kappa_{kz}v_{i}^{[\ell]}v_{k}^{[\ell]}}{\sqrt{\left(\lambda^{[\ell]}-m\omega_{c}^{2}\right)^{2}+\alpha^{2}\omega_{c}^{2}}},$ (23)
$\sin\delta^{[\ell]}(\omega_{c})$$\coloneqq$ $-\frac{\alpha\omega_{c}}{\sqrt{\left(\lambda^{[\ell]}-m\omega_{c}^{2}\right)^{2}+\alpha^{2}\omega_{c}^{2}}}.$ (24)
Here, the superscript (k) indicates that the perturbation caused by $\varphi_{z}$ enters $S$ at node $k$. For each mode $\ell$, the nodal amplitudes depend on the mismatch between the driving frequency $\omega_{c}$ and the eigenmodes of the network associated with the $\lambda^{[\ell]}$, as well as the localization of the corresponding eigenvectors at nodes $i$ and $k$. The damping slightly shifts the location of the resonance peaks in the amplitude away from $\omega_{c}=\pm\sqrt{\lambda^{[\ell]}/m}$ and makes their height finite.
Combining the modal responses linearly provides us with a trajectory for the $\vartheta_{i}$,
$\kappa_{kz}\vartheta_{i}^{(k)}(t)=\sum_{\ell=1}^{N-1}a_{i}^{(k)[\ell]}\sin\left(\omega_{c}t+\delta^{[\ell]}\right).$ (25)
From the trajectory of $\vartheta_{k}$ we can determine the time averaged energy flow $p_{s}=\langle p_{zk}\rangle$ in terms of the amplitudes and phase lags:
$p_{s}(\omega_{c})$ $=$ $-\frac{\kappa_{kz}}{2}\sum_{\ell=1}^{N}a_{k}^{(k)[\ell]}(\omega_{c})\sin\delta^{[\ell]}(\omega_{c})$ (26)
$=$ $\frac{\kappa_{zk}^{2}}{2}\sum_{\ell=1}^{N-1}\frac{\alpha\omega_{c}\left(v_{k}^{[\ell]}\right)^{2}}{\left(\lambda^{[\ell]}-m\omega_{c}^{2}\right)^{2}+\alpha^{2}\omega_{c}^{2}},$ (27)
which is the exact same as in Eq. (21). In fact, the terms in the denominators in the sum over all modes in Eq. (21) resemble the resonance curve of linear oscillators. It is known as Cauchy–Lorentz distribution, Lorentz(ian) function, or Breit–Wigner distribution. Due to the generalization to the resonance of a network $S$, the Lorentzians are summed over all $N-1$ network modes associated to the $\lambda^{[\ell]}$. Each contribution is weighted by the localization of the corresponding eigenvector $\vec{v}^{[\ell]}$ at node $k$, quantified by $\left(v_{k}^{[\ell]}\right)^{2}$. This is how the topology affects the distribution of resonance in the network, see also Zhang et al. (2019).
@@
We remark that both approaches used to derive Eq. (21) straightforwardly generalize to a solitary node connected to several nodes in $S$, see SM I.D. The back-reactions of the network combine linearly, and we simply sum over all neighbors of $z$ (Eq. (S74)).
Furthermore, a treatment of heterogeneous system parameters within $S$ is also possible, see SM I.E.
VI Evaluation
Here, we make some general statements about the properties and solutions of our main result, Eq. (21), and provide some examples.
First, note that the obtained expression for $p_{s}=\langle p_{zk}\rangle$ in Eq. (21), has the same sign as $\omega_{c}$. Therefore, $p_{s}$ is antisymmetric in $\omega_{c}$ and hence in $\omega_{s}\sim\omega_{c}$, cf. Eq. (22). It follows that Eq. (21) is invariant under $\omega_{s}\to-\omega_{s}$ while $P_{z}\to-P_{z}$, i.e., generators and consumers behave similarly with their solitary frequencies having opposite signs. In fact, it is easy to see from Eq. (21) that all solutions lie between zero and the natural frequency, $0\leq|\omega_{s}|\leq|\Omega_{z}|=|P_{z}|/\alpha_{z}$. We see these bounds and symmetry properties confirmed in Fig. 2, and the examples below.
As discussed above, the linear stability of solutions of $Z(\omega_{s})=0$ can be heuristically determined by evaluating the derivative: the orbit at $\omega_{s}$ is indicated to be linearly stable if $\partial_{\omega_{s}}Z(\omega_{s})<0$ and linearly unstable if $\partial_{\omega_{s}}Z(\omega_{s})>0$. This consideration gains importance with the numerical observation that the instantaneous solitary frequency $\dot{\varphi}_{z}(t)$ oscillates with time around $\omega_{s}$, cf. Fig. 1. The solution to $Z(\omega_{s})=0$ is an intersection of $p_{s}(\omega_{s})$ and the straight line $P_{z}-\alpha_{z}\omega_{s}$ with slope $-\alpha_{z}$. Therefore, stability is given if $-\alpha_{z}<\partial_{\omega_{s}}p_{s}(\omega_{s})$. Graphically, this means that $p_{s}(\omega_{s})$ intersects the straight line from below, counting in positive $\omega_{s}$ direction.
The existence of solitary frequencies $\omega_{s}$ as solutions to Eq. (21) relies on sufficiently high eigenvector localization at the neighboring node $k$, quantified by $\left(v_{k}^{[\ell]}\right)^{2}$. Such high localization is present in random networks of high degree heterogeneity Hata and Nakao (2017). Indeed, we observe in Fig. 2 and in the examples below that resonant solitary states are most common at degree-one nodes with a high degree neighbor Nitzbon et al. (2017), and that the solitary frequency scales with the degree $d_{k}$ of node $k$. For an estimate of the relation between the eigenvector localization, the corresponding eigenvalue and the node’s degree, we refer to SM I.A.
The existence of solutions to Eq. (21) also depends on system parameters and their topological distribution. Whether and which solitary states exist in a complex system is a complex question with no easy answer, but we can provide some insights (see SM II.D for more details). Generally, the parameters need to be chosen in an intermediate regime, such that the synchronous state is not globally attractive, but the coupling is also not too weak. Relatively small damping $\alpha_{i}$ and strong coupling $\kappa_{ij}$, such as in power grids, are an indicator for this. We remark that solving the self-consistent Eq. (21) is an efficient way to determine which solitary states exist. Numerical simulations of the dynamics as in Fig. 2 are more precise but more expensive.
@@ VI Evaluation
We now illustrate the power of Eq. (21) using two examples. First we give a minimal effective model that exhibits the resonant tuning mechanism cleanly; then a full complex network with a complex resonant response. We compare the results with numerical simulations Plietzsch et al. (2022). Further examples can be found in the SM II.B.
@@
Figure 3: Network modes shape the landscape of solitary attractor states for the two-node model (inset). (left) Intersections between the straight line $P_{1}-\alpha\omega_{s}$ and the mean energy flow $p_{s}(n=6)$ generate stable and unstable solutions according to Eq. (21). We use $P_{1}=1$, $\alpha=0.1$. The single peak of $p_{s}$ is centered around the network mode close to $\sqrt{Kn}$. (right) Bifurcation diagram with the bifurcation parameter $n$. The two stable branches are connected by an unstable branch and centered around the network mode. Numerical simulations Niehues (2023) confirm predictions of both the location and stability of solitary states.
VI.1 Example I: Two-node model
@@
A minimal model that still shows tristability with two distinct solitary frequencies is a system that neglects the response of the network beyond the node $k$. We have a solitary node ($z=1$), its neighbor ($k=2$), and the rest of the network is assumed to have infinite inertia, and thus no dynamics. This model is shown in the inset in Fig. 3 (right). The neighbor has $n$ links into the infinitely inert part, each with coupling strength $K$. See SM II.A for details.
@@
The left panel in Fig. 3 shows the graphical solution of Eq. (21) for this example system. We see the energy absorbed in node 1 by frequency adaptation, a straight line given by $P_{1}-\alpha\omega_{s}$, and the energy flow due to the network’s response, $p_{s}(\omega_{s})=\langle p_{12}\rangle(\omega_{s})$. The energy flow has a single peak centered around the only network mode, which is close to $\lambda\approx Kn$. At the intersections of these two curves, we have zeros of Eq. (21). However, the linear stability heuristic suggests that only intersections with $\partial_{\omega_{s}}Z<0$ are candidates for stable solitary states. Indeed, numerical simulations confirm the existence of the stable solutions, and that they correspond to the intersections, where $p_{s}$ comes from below.
@@
Repeating this solution process for Eq. (21) for a range of $n$, we can draw the bifurcation diagram in Fig. 3 (right). Each vertical slice corresponds to the top view of an intersection plot like the one in the left panel. For this minimal example system, we can obtain an explicit expression for several branches $n(\omega_{s})$ of stable or unstable solutions (see SM II.A) that correspond to the intersections in the left panel. These branches predict that for $n\lesssim 12$, there is a pair of a stable and an unstable solution $\omega_{s}$ due to the resonance peak around $\lambda$, and a third, stable solution at $\omega_{s}\approx\Omega_{1}=P_{1}/\alpha$. For larger values of $n$, the second and third solutions annihilate, while the first converges to $\omega_{s}\to\Omega_{1}$. Numerical simulations with random initial conditions Niehues (2023) closely correspond to the predicted frequency and stability of solitary states, confirming the existence of tristability between the synchronous state and solitary states at multiple frequencies, and confirming the overall accuracy of our results.
In summary, this minimal example system shows that the resonant (intermediate) solitary state and the natural (decoupled) solitary state are two ends of a spectrum of frequencies, and connected by an unstable branch of solutions. Further, the natural frequency is an upper bound for solitary frequencies and can be approached by tuning the resonant network mode higher.
VI.2 Example II: Complex network